# Exploring the Interior Structure and Mode of Tidal Heating in Enceladus

### v3

@@ Appendix F Effect of shell heterogeneous structure on the Love numbers
Here, we assess the effect of lateral variations in ice shell thickness on the tidal response and its effect on the required measurements of $k_{2m}$ and $h_{2m}$. We use the shape model shown in Figure 11.We build on the analysis of Berne et al. (2023a) and conduct simulations of tidal deformation for four models: 1. without lateral variations in ice shell thickness; 2. with lateral variations in ice shell thickness; 3. with a dissipative core and without lateral variations in ice shell thickness; and 4. with a dissipative core and lateral variations in ice shell thickness. For each model class, we vary the mean ice shell thickness across 15–50 km consistent with the findings in Section 6.1. We assume a core of viscosity $10^{20}$ Pa.s for the case of a non-dissipative core, and $10^{12}$ Pa.s for the case of a dissipative core. All other parameters are chosen to be the same as those used in Figures 1 and 2. To compute $k_{2m}$ and $h_{2m}$, we use a numerical finite element model (Berne et al., 2023a).
Figure 19: Top: Gravitational Love numbers, $k_{20}$ (left column) and $k_{22}$ (right column), against mean ice shell thickness for models that assume an ice shell of constant thickness (dotted line), an ice shell with lateral variations in thickness (solid line), or a dissipative core (dotted and solid red lines). Bottom: Similar to top row except showing displacement Love numbers $h_{20}$ and $h_{22}$ plotted against mean ice shell thickness. x-axes are plotted on a log-scale.

### v5

@@ Appendix F Effect of shell heterogeneous structure on the Love numbers
Here, we assess the effect of lateral variations in ice shell thickness on the tidal response and its effect on the required measurements of $k_{2m}$ and $h_{2m}$. We use the shape model shown in Figure 11. We build on the analysis of Berne et al. (2023a) and conduct simulations of tidal deformation for four models: 1. without lateral variations in ice shell thickness; 2. with lateral variations in ice shell thickness; 3. with a dissipative core and without lateral variations in ice shell thickness; and 4. with a dissipative core and lateral variations in ice shell thickness. For each model class, we vary the mean ice shell thickness across 15–50 km consistent with the findings in Section 5.2. We assume a core of viscosity $10^{20}$ Pa.s for the case of a non-dissipative core, and $10^{12}$ Pa.s for the case of a dissipative core. All other parameters are chosen to be the same as those used in Figures 2 and 3. To compute $k_{2m}$ and $h_{2m}$, we use a numerical finite element model (Berne et al., 2023a).
Figure 22: Top: $k_{20}$ (left column) and $k_{22}$ (right column), against mean ice shell thickness for models that assume an ice shell of constant thickness (dotted line), an ice shell with lateral variations in thickness (solid line), or a dissipative core (dotted and solid red lines). Bottom: Similar to top row except showing displacement Love numbers $h_{20}$ and $h_{22}$ plotted against mean ice shell thickness. x-axes are plotted on a log-scale.